OpenStudy (safyousuff97):

Given the function f(x) = 2(3)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

OpenStudy (safyousuff97):

@baru

OpenStudy (baru):

average rate of change= change in f(x) / change in x

OpenStudy (baru):

lets find the average rate of change for section A. if we go from x=0 to x=1, how much has x changed?

OpenStudy (safyousuff97):

1

OpenStudy (baru):

yes, so 1 is the denominator in the formula for average rate of change now we need to find how much f(x) changes when we go from x=0 to x=1 to do that find f(x) at x=0 and find f(x) at x=1 what do you get?

OpenStudy (safyousuff97):

i'm not sure

OpenStudy (baru):

just substitute x=0 and x=1 in f(x)

OpenStudy (baru):

when x=0 f(x)=2(3)0=?

OpenStudy (baru):

are you sure you have written the quesiton correctly? f(x)=2(3)x specifically the (3) part?

OpenStudy (safyousuff97):

i mean \[f(3)^{x}\]

OpenStudy (safyousuff97):

wait one sec

OpenStudy (baru):

i dont get you... can you type the whole question

OpenStudy (safyousuff97):

\[f(x)=2(3)^{x}\]

OpenStudy (baru):

ok, can you find f(x) when x=0?

OpenStudy (safyousuff97):

no

OpenStudy (safyousuff97):

coz it 0

OpenStudy (baru):

any number raised to the power 0, is 1

OpenStudy (baru):

can you find f(x) when x=0 now?

OpenStudy (safyousuff97):

yes bcoz it will be 1 no right?

OpenStudy (safyousuff97):

*now

OpenStudy (baru):

yes, so f(x)=?

OpenStudy (baru):

\[f(x)=2(3)^{0}=?\]

OpenStudy (safyousuff97):

3

OpenStudy (baru):

\(3^0\)=1

OpenStudy (baru):

how did you get 3?